The suspension system of a 1900 kg automobile "sags" 13 cm when the chassis is placed on it. Also, the oscillation amplitude decreases by 52% each cycle. Estimate the values of (a) the spring constant k and (b) the damping constant b for the spring and shock absorber system of one wheel, assuming each wheel supports 475 kg.
Answer #1

a)

From Hooke's law

K=mg/x =(475*9.8/0.13)

K=35807.7 N/m

b)

Since period of shm

T=2pi*sqrt(m/K) =2pi*sqrt(475/35807.7) =0.723667 s

From

X=XMaxe-bT/2m

Given X=(1-0.52)Xmax

0.48= e-b*0.723667/2*475

Ln(0.48)=-b*(7.6175*10-4)

b=963.52 Kg/s